OpenStudy (anonymous):

What are the explicit equation and domain for a geometric sequence with the first term of 5 and a second term of -10?

OpenStudy (anonymous):

an=5(-2)^(n-1); all integers when n≥1 an=5(-2)^(n-1); all integers where n≥0 an=5(-15)^(n-1); all integers where n≥1 an=5(-15)^(n-1); all integers where n≥0

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

yay @phi is here

OpenStudy (anonymous):

I'm pretty sure the second half is n≥1

OpenStudy (anonymous):

right?

OpenStudy (phi):

You could test each choice. You should get 5, -10 as the first two numbers.

OpenStudy (anonymous):

a

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

b

OpenStudy (anonymous):

no a

OpenStudy (anonymous):

omg im so confused

OpenStudy (phi):

***an=5(-2)^(n-1); all integers when n≥1*** for n=1 you get \[ a_1 = 5(-2)^{1-1} \\ a_1 = 5\cdot (-2)^0 \] anything to the zero power is 1 , so that simplifies to \[ a_1= 5 \cdot 1 \\a_1=5\] that looks good so far

OpenStudy (phi):

now find a2 for choice A, using n=2

OpenStudy (anonymous):

a2=5(-2)^(2-1) =5(-2)^1 =-10

OpenStudy (phi):

yes. so choice A looks like the answer. You could double check the other three, but they won't work (give 5, -10)

OpenStudy (anonymous):

ok thank you i have another but i don't know how to post the graphs so oh well

OpenStudy (phi):

do you know how to take a screen shot?

OpenStudy (anonymous):

the wording is horrible

OpenStudy (anonymous):

yeah put its on my other computer hang on

OpenStudy (anonymous):

It says which logarithmic graph can be used to approximate the value of y in the equation 2^y=3

OpenStudy (anonymous):

i don't know how to log that?

OpenStudy (phi):

take the log of both sides. Log base 2 is the most convenient, but log base 10 or log base 3 (natural log) all can be used. use the "rule" \[ \log(a^b) = b \log(a) \]

OpenStudy (anonymous):

so… log10 2^y= log10 3

OpenStudy (phi):

and then use the rule to write \[ y \log 2 = \log 3 \\ y = \frac{\log 3}{\log 2} \]

OpenStudy (phi):

to get the answer from a graph, we need to figure out what log base they used. for example, what is the x value that corresponds to y=1 (that should give us the base)

OpenStudy (phi):

Here are 3 possible logs if they gave you log base 2, read off the y value at x=3

OpenStudy (anonymous):

thank you i got it